QUESTION 2 Given: f(x) = – 2x²+x-5 Find : f(-6)= Of(-6) = -83 f(-6) = 133 f(-6) = 25 None of the above
Answer #1

: Given, ifa --28+82-5 f66) 2-266)²+(-6.)-5 .f(-5) = -2(36) -6-5 f66) = -720-11 |f66)3-83) The auswer i first option.